OpenStudy (anonymous):

needing help with my fractions that i have to make improper, mixed numbers and lowest rates

OpenStudy (texaschic101):

you got any questions ?

OpenStudy (anonymous):

yes i do and i just can't get it down

OpenStudy (texaschic101):

You just can't get it down ? you can't understand it, or you just can't post the question.

OpenStudy (anonymous):

both

OpenStudy (anonymous):

ok just so that you know what i am doing i am in my GED book studying the decimals and fractions

OpenStudy (texaschic101):

do you know how to turn mixed numbers into improper fractions ?

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

thats the hard part because i never had a math class that taught me this so now that i am getting my GED i dont understand it

OpenStudy (texaschic101):

Let us begin..lol 2 1/4 -- take the whole number (2) and multiply it by the denominator (4), then add that to the numerator (1), then put the whole thing over the original denominator. Thus giving us : 9/4 (an improper fraction) Those are the basic steps for turning a mixed number into an improper fraction. 1 2/3 = (1 x 3 + 2) / 3 = 5/3 3 3/8 = (3 x 8 + 3) / 8 = 27/8 starting to understand ?

OpenStudy (anonymous):

im understanding...yes

OpenStudy (texaschic101):

if you didn't know...an improper fraction is when the numerator is larger then the denominator

OpenStudy (texaschic101):

what about adding,subtracting,multiplying,dividing fractions ? Do you know how to do those ? Do you need to know ?

OpenStudy (anonymous):

so for the ones that i have in my book are adding and subtracting

OpenStudy (texaschic101):

One very important rule about adding and subtracting fractions...it can only be done if they have the same denominator.

OpenStudy (texaschic101):

Let me ask you this.. are you familiar with equivalent fractions ?

OpenStudy (anonymous):

 i seems to keep getting the wrong answer for these and i need help finding a strategy to hep me solve it . like steps that i can remember